INTEGRATION OF TWISTED DIRAC BRACKETS

HENRIQUE BURSZTYN MARIUS CRAINIC ALAN WEINSTEIN CHENCHANG ZHU

Abstract. Given a Lie groupoid G over a manifold M, we show that multiplicative 2-forms on G relatively closed with respect to a closed 3-form ϕ on M correspond to maps from the Lie algebroid of G into T^{∗}M satisfying an algebraic condition and a differential condition with respect to the ϕ-twisted Courant bracket. This correspondence describes, as a special case, the global objects associated to ϕ-twisted Dirac structures. As applications, we relate our results to equivariant cohomology and foliation theory, and we give a new description of quasi-hamiltonian spaces and group-valued momentum maps.